OpenStudy (anonymous):

what is the period and asymptote of the graph tan3(x+2Pi/3)?

OpenStudy (anonymous):

what is tan3 you mean tan^3

OpenStudy (anonymous):

Tan times 3

OpenStudy (anonymous):

\[\tan^3(x+\frac{\pi}{3})\]

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

What is the period and asymptote?

OpenStudy (anonymous):

\[3\tan(x+\frac{\pi}{3})\]

OpenStudy (anonymous):

no maths mind its tan 3

OpenStudy (anonymous):

What is the period and asymptote?

OpenStudy (zpupster):

period of tan is pi(180 deg.) = tan(bx) = pi/b or 180/3 = 60

OpenStudy (zpupster):

i beleive the va 2pi/3. and period is pi/3